20,000 raffle tickets are sold one first prize of $1000 two second prices of $500 and three thied prizes of $100 each will be awarded with all winners selected randomly. if you purchased one ticket what are your expected growth winnings .
Answer #1

Given That

raffle tickets =20,000

probability of winning first prize = p1 = 1/20000

probability of winning a 2nd prize = p2 = 2/20000 = 1/10000

probability of winning a 3rd prize = p3 = 20/20000 = 1/1000

E = expectation value = Sum_i=1..3 pi * Vi, Vi = value of prize, pi = probability of winning prize,

E = 1/20000*1000 + 1/10000*300 + 1/1000*10 = 1/20 + 3/100 + 1/100 = 9/100 = 9 cents = expected winnings for a person buying one ticket.